OpenStudy (anonymous):

The length of a rectangle is 2 inches longer than its width. If the perimeter is 36 inches, find the length and width

OpenStudy (yanasidlinskiy):

\(\Huge\bf \color{yellow}{Welcome~to~OpenStudy!!}\hspace{-310pt}\color{cyan}{Welcome~to~OpenStudy!!}\hspace{-307.1pt}\color{midnightblue}{Welcome~to~\color{purple}{Open}}\color{blue}{Study!!!!}\) The perimeter of a rectangle is \(2 \times (length + width)\)

OpenStudy (yanasidlinskiy):

Got that so far?

OpenStudy (anonymous):

Yes

OpenStudy (yanasidlinskiy):

We can take width as x, then length is 4 inches more, so: Width = x Length = x + 4 Equation we get: \(⇒2(x+x+4)=36\) Next step would be to?

OpenStudy (anonymous):

Would the equation be x+x+x+2+x+2

OpenStudy (yanasidlinskiy):

No. The next step would be to divide by 2. After we have divided by 2, Our equation now looks like this: x+x+4=18 Do you know what the next step might be?

OpenStudy (shadowlegendx.):

@shamil98